# MFTCXplain: A Multilingual Benchmark Dataset for Evaluating the Moral Reasoning of LLMs through Hate Speech Multi-hop Explanations

### v3

@@ Abstract
Ensuring the moral reasoning capabilities of Large Language Models (LLMs) is a growing concern as these systems are used in socially sensitive tasks. Nevertheless, current evaluation benchmarks present two major shortcomings: a lack of annotations that justify moral classifications, which limits transparency and interpretability; and a predominant focus on English, which constrains the assessment of moral reasoning across diverse cultural settings. In this paper, we introduce MFTCXplain, a multilingual benchmark dataset for evaluating the moral reasoning of LLMs via hate speech multi-hop explanation using Moral Foundation Theory (MFT). The dataset comprises 3,000 tweets across Portuguese, Italian, Persian, and English, annotated with binary hate speech labels, moral categories, and text span-level rationales. Empirical results highlight a misalignment between LLM outputs and human annotations in moral reasoning tasks. While LLMs perform well in hate speech detection (F1 up to 0.836), their ability to predict moral sentiments is notably weak (F1 < 0.35). Furthermore, rationale alignment remains limited mainly in underrepresented languages. These findings show the limited capacity of current LLMs to internalize and reflect human moral reasoning.
00footnotetext: Warning: This document contains offensive content.
@@ 1 Introduction
LLMs perform tasks that rely on moral reasoning, from moderating harmful content to providing an ethical advisor (Dillion et al., 2025). However, recent research has raised concerns about the robustness, explainability, and alignment of LLMs with diverse human values (Oh and Demberg, 2025; Huang et al., 2025; Agarwal et al., 2024; Atari et al., 2023b; Abdurahman et al., 2024), as they often exhibit inconsistent behavior, lack transparency, and reflect culturally contingent moral frameworks.
@@
Figure 1: Hate speech multi-hop explanation.
@@
One domain where moral reasoning is both prominent and consequential is in hate speech. People often use moral language around purity, loyalty, and authority, to justify incivility or portray harm to the outside world as righteous (Kennedy et al., 2023; Fiske and Rai, 2014). These moral framings of hate trigger moral outrage, increase engagement, and reinforce group loyalty (Brady et al., 2020; Abdurahman et al., 2025). Measuring this type of moral reasoning is crucial, as different moral appeals shape hate in culturally and psychologically distinct ways (Graham et al., 2013; Hoover et al., 2021). Despite advances in hate speech detection in terms of linguistic and social bias (Davani et al., 2023; Plaza-del arco et al., 2023; Vargas et al., 2021; Davidson et al., 2019), models often remain “black-box”, producing classifications without revealing the underlying reasoning (Kennedy et al., 2020). In addition, most existing benchmarks are English-only, constraining the development of models capable of culturally grounded moral reasoning (Buscemi et al., 2025).
@@
To address these challenges, we introduce MFTCXplain [^1] The dataset and code, and results are available at https://github.com/franciellevargas/MFTCXplain, (Moral Foundations Twitter Corpus with EXplanations), the first multilingual benchmark dataset for evaluating moral reasoning in LLMs using expert-annotated rationales based on Moral Foundations Theory (Graham et al., 2013). The dataset comprises 3,000 tweets in four languages—Portuguese, Persian, Italian, and English—annotated for hate speech, fine-grained moral categories, and text spans (rationales) that justify each moral label. Notably, the Persian and Brazilian Portuguese portions represent the Global South, low-resource contexts often neglected in NLP research, while Italian, though not from the Global South, remains understudied. Annotations were conducted by native speakers with diverse cultural backgrounds, and detailed annotator metadata is provided to support bias analysis (Prabhakaran et al., 2021; Davani et al., 2024).
@@
While multi-hop explanation has traditionally been studied in question answering (Jhamtani and Clark, 2020; Valentino et al., 2021; Thayaparan et al., ) and fact verification (Ma et al., 2024), we draw inspiration from this paradigm to frame moral explanation as a multi-hop reasoning process, as shown in Figure 1. This requires LLMs to connect expressions of hate with moral principles through intermediate inferences, such as identifying hate speech (offensive content against targeted groups) - hop 1; recognizing and mapping moral violations - hop 2; and their rationales - hop 3; thus requiring multi-level moral reasoning and compositional inference. To assess current model capabilities, we benchmark state-of-the-art LLMs including GPT-4o and LLaMA 3.3 70B using zero-shot, few-shot, and Chain-of-Thought (CoT) prompting. Our results reveal substantial divergence between human and model-generated explanations and low macro F1 scores (often below 0.35) when predicting moral sentiment, highlighting the challenges LLMs face in aligning with culturally human moral reasoning. The dataset, including annotation agreements/disagreements, model outputs, and code, is publicly available to facilitate future research.
@@
Our contributions are summarized as follows: (i) We provide the first expert-annotated multilingual dataset of hate speech multi-hop explanation for evaluating LLMs’ moral reasoning; (ii) We propose a novel annotation schema based on Moral Foundations Theory to enhance the interpretability of hate speech explanations by multi-hop explanation; (iii) We conduct the first multilingual study on explainable moral reasoning in LLMs, revealing insights across cultures.
@@ 2 Related Work
Authors Datasets Languages # Instances Platforms Models F-score
Johnson and Goldwasser (2018a) Non-Name EN 2,050 Twitter SVM 0.72
Hoover et al. (2020a) MFTC EN 35,108 Twitter SVM, LSTM 0.80
@@
Beiró et al. (2023) Non-Name EN 4,811 Facebook LSTM, Regression 0.82
Table 1: Overview of annotated datasets with Moral Foundation Theory (MFT).
Authors Datasets Languages # Instances Platforms Models F-score
@@
Pavlopoulos et al. (2021) Non-Name EN 10,629 Civil Comments dataset BERT 0.71
@@
Salles et al. (2025) HateBRXplain PT 7,000 Instagram BERT, DistilBERT, PTT5 0.91
@@
2.1 Moral Foundation Theory Annotated Datasets
Previous works have created datasets annotated with MFT (See Table 1). The Moral Foundations Twitter Corpus (MFTC) Hoover et al. (2020b) includes over 35,000 tweets annotated with the five MFT pairs proposed by Graham et al. (2013): care/harm, fairness/cheating, loyalty/betrayal, authority/subversion, and purity/degradation. Similarly, the Moral Foundations Reddit Corpus (MFRC) Trager et al. (2022) contains approximately 16,000 Reddit comments annotated across eight moral dimensions based on the updated MFT proposed by Atari et al. (2023a).Johnson and Goldwasser (2018b) introduced a dataset of over 2,000 politician tweets labeled for the original five MFT pairs. Additional work during the COVID-19 pandemic extended moral analysis Beiró et al. (2023), examining 500,000 Facebook posts through the lens of Liberty, Care, and Authority, while Pacheco et al. (2022) analyzed 750 annotated tweets linking vaccine attitudes to the five MFT dimensions (Haidt and Graham, 2007), including Liberty/Oppression. While these datasets have enriched our understanding of moral language, they lack human-annotated rationales that explain why a specific moral label applies, limiting their ability for training interpretable models.
@@ 2.2 Hate Speech Datasets with Human-Annotated Rationales
In contrast, several hate speech datasets include human-annotated rationales to improve transparency and model explainability (See Table 2). Pioneering work by Zaidan and Eisner (2008) and the ERASER benchmark (DeYoung et al., 2019) demonstrate the value of such rationales. In the hate speech domain, HateXplain (Mathew et al., 2021) includes labels, targets, and rationale spans, revealing improvements in performance and bias mitigation. Other rationale-based resources have expanded into low-resource and multilingual contexts, including DOSA for Tamil-English and Kannada-English (Ravikiran and Annamalai, 2021), ViHOS for Vietnamese (Hoang et al., 2023), HateBRXplain for Brazilian Portuguese (Salles et al., 2025), and PHATE for Persian (Delbari et al., 2024a). While these datasets improve explainability, they generally lack moral value annotations, limiting deeper of hate speech across cultures.
3 MFTCXplain Corpus
3.1 Data Collection
We collected 3,000 tweets across four languages with 704, 621, 608 and 1,067 samples for English, Italian, Persian, and Portuguese, respectively. These documents were extracted from a variety of benchmark hate speech datasets, as shown in Table 3. Tweets were selected based on relevance, annotation availability, and alignment with our hate speech definition, with some re-annotated for consistency across languages. We defined hate speech as a type of offensive language used against groups based on their social identity (Fortuna and Nunes, 2018; Zampieri et al., 2019; Vargas et al., 2022; Wardle, 2024). A complete description of dataset sources and annotation procedures is also included in the Appendix A.
@@ 3.2 Annotation Process
Each tweet in MFTCXplain was labeled for 10 categories of moral sentiment as described in an adapted version of the Moral Foundations Coding Guide (See Appendix A). First, annotators decided whether the tweet contained any kind of moral sentiment or was labeled as Non-Moral. If labeled as containing moral sentiment, annotators were then asked to label at least 1 and at most 3 categories of moral sentiment present in the tweet, in order of the most salient moral value to the least. Finally, annotators were asked to highlight the rationale or the span of the text that guided each of their label(s). We also gathered annotator profile meta-data (e.g, psycho-cultural demographics) to facilitate bias analysis (See Appendix A).
@@ 3.2.1 Moral Categories
For annotating moral sentiment, our annotation schema is based on Moral Foundations Theory (Graham et al., 2013), which identifies five core domains of moral reasoning believed to be universal across cultures (Atari et al., 2023a). Each domain is framed as a bipolar dimension, with a virtue (moral adherence) and a vice (moral violation) representing contrasting ends of the same moral spectrum. This framework has been widely used in prior research in computer science (e.g., Hoover et al. (2020a)) and forms the basis of the moral annotation for hate speech multi-hop explanations in MFTCXplain corpus. The five foundational domains are defined as follows:
@@
Although these foundations are often conceptualized as binary dimensions, prior research has shown that virtue and vice expressions in natural language are often independent (Hoover et al., 2020a). For instance, a tweet condemning harm may not necessarily express care. To account for this, our annotations treat virtues and vices as separate categories, allowing for more precise and nuanced labeling of moral content
3.2.2 Moral Rationales
Speech Type Tweet (with highlighted rationale spans) Moral
@@ 3.2.3 Annotation Evaluation
Rationales Jaccard F1 BERTScore
Rationale 1 0.590 0.690 0.886
Rationale 2 0.632 0.700 0.906
Rationale 3 0.551 0.626 0.909
Table 5: Jaccard, F1 and BERTScore for rationales evaluation of human annotators for the Portuguese corpus.
@@ 4 Corpus Analysis
Morality of Hate Speech Across Cultures: We also analyzed the moral framing of hate speech across languages, finding that tweets labeled as hate speech rarely lacked moral content, as shown in Appendix A, Figure 7. Instead, they overwhelmingly relied on moral violations, especially Harm, Degradation, and Betrayal, consistent with prior work on the moral rhetoric of hate (Kennedy et al., 2023). While vice framing was dominant, some hate speech tweets drew on virtues such as Loyalty or Fairness to morally justify outgroup hostility. These trends were largely consistent across English, Italian, Portuguese, and Persian, though the relative prominence of each foundation varied. For example, Portuguese and Persian tweets more often presented fairness violations, whereas Italian and English emphasized degradation. These patterns suggest that moral appeals are central to the way hate is expressed online, but their specific form is shaped by the cultural context. For a more in-depth analysis, see Appendix C.
5 Experimental Setup
@@
To measure how state-of-the-art LLMs align with the moral values and hate speech labels annotated by our experts, we utilized the technique of LLMs as objective judges (Zheng et al., 2023; Verga et al., 2024), and extended it with a multi-hop explanation evaluation strategy. This approach requires models not only to predict the labels of the input text, but also to infer the corresponding moral violation, with a brief evidence-based justification, specifying the rationales or group of words that triggered the moral value classification. We selected LLaMA-70B-Instruct [^3] https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct and GPT-4o Mini  [^4] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/, accessing them all via API. The prompting experiments were conducted on Google Colab using the CPU for processing, with the Macro F1 score metric from Sklearn library. LLaMA-70B-Instruct was chosen for its established effectiveness in similar evaluations Piot and Parapar (2025) while GPT-4o-Mini provides an excellent balance of performance and cost-efficiency. Moreover, using both models enables for a comparative analysis between an open (LLaMA) and closed-source model (GPT-4o-Mini) for our research tasks. Parameter settings remained consistent across both models, with each being tasked with determining: (i) Whether the tweet comprises hate speech; (ii) Which moral sentiments are present in the tweet; (iii) Offering rationales for the moral sentiment prediction via the relevant text spans extracted from the tweet.
@@
To obtain comprehensive assessments, we implemented three distinct prompting strategies: (i) zero-shot prompting with no examples provided, asking the model to make determinations based solely on its knowledge; (ii) few-shot prompting with four examples included to help guide the model behavior; and Chain-of-Thought (CoT) strategy, with a step-by-step reasoning approach aimed at guiding the models’ reasoning. The examples of prompts are shown in the Appendix D.

### v4

# MFTCXplain: A Multilingual Benchmark Dataset for Evaluating the Moral Reasoning of LLMs through Multi-hop Hate Speech Explanation

@@ Abstract
Ensuring the moral reasoning capabilities of Large Language Models (LLMs) is a growing concern as these systems are used in socially sensitive tasks. Nevertheless, current evaluation benchmarks present two major shortcomings: a lack of annotations that justify moral classifications, which limits transparency and interpretability; and a predominant focus on English, which constrains the assessment of moral reasoning across diverse cultural settings. In this paper, we introduce MFTCXplain, a multilingual benchmark dataset for evaluating the moral reasoning of LLMs via multi-hop hate speech explanations using the Moral Foundations Theory. MFTCXplain [^1] MFTCXplain dataset is available at https://github.com/franciellevargas/MFTCXplain. comprises 3,000 tweets across Portuguese, Italian, Persian, and English, annotated with binary hate speech labels, moral categories, and text span-level rationales. Our results show a misalignment between LLM outputs and human annotations in moral reasoning tasks. While LLMs perform well in hate speech detection (F1 up to 0.836), their ability to predict moral sentiments is notably weak (F1 < 0.35). Furthermore, rationale alignment remains limited mainly in underrepresented languages. Our findings show the limited capacity of current LLMs to internalize and reflect human moral reasoning [^2] Warning: This document contains offensive content..
@@ 1 Introduction
Large Language Models (LLMs) perform tasks that rely on moral reasoning, from moderating harmful content to providing an ethical advisor (Dillion et al., 2025). However, recent research has raised concerns about the robustness, explainability, and alignment of LLMs with diverse human values (Oh and Demberg, 2025; Huang et al., 2025; Agarwal et al., 2024; Atari et al., 2023b; Abdurahman et al., 2024), as they often exhibit inconsistent behavior, lack transparency, and reflect culturally contingent moral frameworks.
@@
Figure 1: Multi-hop hate speech explanation for moral reasoning evaluation.
@@
One domain where moral reasoning is especially prominent is hate speech. People often use moral language (e.g., purity, loyalty, and authority) to justify incivility or frame out-group harm as morally righteous (Kennedy et al., 2023; Fiske and Rai, 2014). Such moral framings of hate can trigger moral outrage, increase engagement, and reinforce group loyalty (Brady et al., 2020; Abdurahman et al., 2025; Trager et al., 2025b). Measuring this form of moral reasoning is crucial, as different moral appeals shape hate in culturally and psychologically distinct ways (Graham et al., 2013; Hoover et al., 2021).
Despite advances in hate speech detection that address linguistic and social bias (Davani et al., 2023; Vargas et al., 2021; Davidson et al., 2019), models often remain “black-box” systems, producing classifications without revealing their underlying reasoning (Kennedy et al., 2020). Furthermore, most existing benchmarks are English-only, limiting the development of models capable of culturally grounded moral reasoning (Buscemi et al., 2025; Plaza-del-Arco et al., 2023).
@@
To address these challenges, we introduce MFTCXplain (Moral Foundations Twitter Corpus with EXplanations), the first multilingual benchmark dataset for evaluating moral reasoning in LLMs using expert-annotated rationales based on the Moral Foundations Theory (MFT) (Graham et al., 2013). The dataset comprises 3,000 tweets in four languages (Portuguese, Persian, Italian, and English) annotated for hate speech, fine-grained moral categories, and text spans (rationales) that justify each moral label. Notably, Persian and Brazilian Portuguese represent the Global South, low-resource contexts often neglected in NLP research, while Italian, though not from the Global South, remains understudied. Annotations have been conducted by native speakers with diverse cultural backgrounds. We provide detailed annotator metadata to support bias analysis.
@@
While multi-hop explanation has traditionally been studied in question answering (Nezhad and Agrawal, 2025; Jhamtani and Clark, 2020; Valentino et al., 2021) and fact verification (Ma et al., 2024), we draw on this paradigm to conceptualize moral explanation as a multi-hop reasoning process (see Figure 1). This requires LLMs to connect expressions of hate with moral sentiments through intermediate inferences, such as identifying hate speech (offensive language used against targeted groups) - hop 1; recognizing moral violations - hop 2; and their rationales - hop 3; thus requiring multi-level moral reasoning and compositional inference.
To evaluate current model capabilities, we benchmark state-of-the-art LLMs, including GPT-4o and LLaMA 3.3 70B, using zero-shot, few-shot, and Chain-of-Thought (CoT) prompting. Our results reveal a substantial divergence between human- and model-generated explanations and low macro F1 scores (often below 0.35) when predicting moral sentiment, highlighting the challenges LLMs face in aligning with culturally human moral reasoning. The dataset, including annotation agreements and disagreements, model output, and code, is publicly available to facilitate future research.
@@
Our contributions are summarized as follows: (i) We provide the first expert-annotated multilingual dataset of multi-hop hate speech explanation for evaluating LLMs’ moral reasoning; (ii) We propose a novel annotation schema based on the MFT to enhance the interpretability of hate speech detection models by multi-hop explanation; (iii) We conduct the first multilingual study on explainable moral reasoning in LLMs, revealing insights across cultures and four languages.
@@ 2 Related Work
Authors Datasets Lang # Instances Platforms Models F-score
Johnson and Goldwasser (2018) No-Name EN 2,050 Twitter SVM 0.72
Hoover et al. (2020) MFTC EN 35,108 Twitter SVM, LSTM 0.80
@@
Beiró et al. (2023) No-Name EN 4,811 Facebook LSTM, Regression 0.82
Trager et al. (2025a) MFTCXplain Multi 3,000 Twitter GPT, LLaMA 0.83
Table 1: Overview of datasets from the literature annotated with Moral Foundations Theory.
Authors Datasets Lang # Instances Platforms Models F-score
@@
Pavlopoulos et al. (2021) No-Name EN 10,629 Civil Comments BERT 0.71
@@
Salles et al. (2025) HateBRXplain PT 7,000 Instagram BERT, DistilBERT 0.91
@@
2.1 Moral Foundations Theory Annotated Datasets
Previous work has created datasets annotated with MFT (see Table 1). The Moral Foundations Twitter Corpus (MFTC) Hoover et al. (2020) includes over 35,000 tweets annotated with the five MFT pairs proposed by Graham et al. (2013): care/harm, fairness/cheating, loyalty/betrayal, authority/subversion, and purity/degradation. Similarly, the Moral Foundations Reddit Corpus (MFRC) Trager et al. (2022) contains approximately 16,000 Reddit comments annotated across eight moral dimensions based on the updated MFT proposed by Atari et al. (2023a). Johnson and Goldwasser (2018) introduced a dataset of over 2,000 politician tweets labeled for the original five MFT pairs. Building and extending on this dataset, Roy et al. (2021) propose a relational learning model to predict moral attitudes towards entities and moral foundations. During the COVID-19 pandemic, for instance, Beiró et al. (2023) analyzed 500,000 Facebook posts in relation to Liberty, Care, and Authority, while Pacheco et al. (2022) examined 750 annotated tweets to link vaccine attitudes with the five MFT dimensions (Haidt and Graham, 2007), including Liberty/Oppression. While these datasets have enriched our understanding of moral language, they lack human-annotated rationales that explain why a specific moral label applies, limiting their ability for studying interpretable models.
2.2 LLMs applied to MFT
Recently, there has been a growing body of research exploring the intersection between LLMs and moral framing. For instance, Zhou et al. (2023) assess LLMs’ understanding and adherence to specific moral theories, focusing on the Dyadic Morality framework Schein and Gray (2018) using COT prompting. While their results show that LLMs demonstrate reasonable understanding of moral reasoning frameworks, they find that no single theory consistently aligns with human annotators. Critically, the authors attribute much of this misalignment to limitations in existing datasets rather than model capabilities, highlighting the need for improved dataset construction and annotation practices. Given that LLMs are trained on vast amounts of online data, Abdulhai et al. (2023) investigate whether moral foundations expressed by LLMs reflect human biases. Their analysis reveals that LLMs tend to emphasize certain moral dimensions over others, and demonstrate that prompt design can significantly influence which moral values the model highlights.
Most relevant to our work is the study by Islam and Goldwasser (2025), which investigates how LLMs can be used to annotate moral framing in online vaccination debates through human-evaluated model explanations. Their findings emphasize the potential of LLMs to enhance annotation accuracy, reduce cognitive load, and simplify complex moral reasoning tasks. However, their study is limited to few-shot prompting, which may not fully capture the breadth of LLM capabilities across more diverse or nuanced moral scenarios. Current studies primarily focus on English-only datasets and use single-prompting approaches without comprehensive evaluation, leaving open questions about how well LLMs handle more diverse and nuanced moral scenarios, and which prompting strategies are most effective across languages.
@@ 2.3 Hate Speech Datasets with Human-Annotated Rationales
Several hate speech datasets include human-annotated rationales to improve transparency and model explainability (see Table 2). Pioneering work by Zaidan and Eisner (2008) and the ERASER benchmark (DeYoung et al., 2019) demonstrates the value of such rationales. In the hate speech domain, HateXplain (Mathew et al., 2021) includes labels, targets, and rationale spans, revealing improvements in performance and bias mitigation. Other rationale-based resources have expanded into low-resource and multilingual contexts, including DOSA for Tamil-English and Kannada-English (Ravikiran and Annamalai, 2021), ViHOS for Vietnamese (Hoang et al., 2023), HateBRXplain for Brazilian Portuguese (Salles et al., 2025), and PHATE for Persian (Delbari et al., 2024a). While these datasets improve explainability, they generally lack moral value annotations, limiting deeper analysis of hate speech across cultures.
@@
3 MFTCXplain Corpus
3.1 Data Collection
We collected 3,000 tweets across four languages with 704, 621, 608 and 1,067 samples for English, Italian, Persian, and Portuguese, respectively. These documents were extracted from a variety of benchmark hate speech datasets, as shown in Table 3.
Tweets were selected based on relevance, annotation availability, and alignment with our hate speech definition, with some re-annotated for consistency across languages. We defined hate speech as a type of offensive language used against groups based on their social identity (Fortuna and Nunes, 2018; Zampieri et al., 2019; Vargas et al., 2022; Wardle, 2024). Dataset sources and annotation procedures are detailed in the Appendix A.
@@ 3.2 Annotation Process
Each tweet in MFTCXplain was labeled for 10 categories of moral sentiment as described in an adapted version of the Moral Foundations Coding Guide (see the Appendix A). First, annotators decided whether the tweet contained any kind of moral sentiment or was labeled as Non-Moral. If labeled as containing moral sentiment, annotators were then asked to label at least 1 and at most 3 categories of moral sentiment present in the tweet, in order of the most salient moral value to the least. Finally, annotators were asked to highlight the rationale or the span of the text that guided each of their label(s). We also gathered annotator profile meta-data (e.g, psycho-cultural demographics) to facilitate bias analysis (see the Appendix A).
@@ 3.2.1 Moral Categories
For annotating moral sentiment, our annotation schema is based on Moral Foundations Theory (Graham et al., 2013), which identifies five core domains of moral reasoning believed to be universal across cultures (Atari et al., 2023a). Each domain is framed as a bipolar dimension, with a virtue (moral adherence) and a vice (moral violation) representing contrasting ends of the same moral spectrum. This framework has been widely used in prior research in computer science (e.g., Hoover et al. (2020)) and forms the basis of the moral annotation for multi-hop hate speech explanations in MFTCXplain corpus. The five foundational domains are defined as follows:
@@
Hate Speech Label(s) Tweet (with highlighted rationale spans) Moral
@@
Although these foundations are often conceptualized as binary dimensions, prior research has shown that virtue and vice expressions in natural language are often independent (Hoover et al., 2020). For instance, a tweet condemning harm may not necessarily express care. To account for this, our annotations treat virtues and vices as separate categories, allowing for more precise and nuanced labeling of moral content
3.2.2 Moral Rationales
@@ 3.2.3 Annotation Evaluation
Rationales Jaccard F1
Rationale 1 0.590 0.690
Rationale 2 0.632 0.700
Rationale 3 0.551 0.626
Table 5: Avg. Jaccard and F1 for rationales evaluation between human annotators of the Portuguese corpus.
@@ 4 Corpus Analysis
Morality of Hate Speech Across Cultures: We also analyzed the moral framing of hate speech across languages, finding that tweets labeled as hate speech rarely lacked moral content, as shown in Appendix A, Figure 7. Instead, they overwhelmingly relied on moral violations, especially Harm, Degradation, and Betrayal, consistent with prior work on the moral rhetoric of hate (Kennedy et al., 2023). While vice framing was dominant, some hate speech tweets drew on virtues as Loyalty or Fairness to morally justify outgroup hostility. These trends were largely consistent across English, Italian, Portuguese, and Persian, though the relative prominence of each foundation varied. For example, Portuguese and Persian tweets often presented fairness violations, whereas Italian and English emphasized degradation. These patterns suggest that moral appeals are central to the way hate is expressed online, but their specific form is shaped by the cultural context (see more in the Appendix C).
@@
To measure how state-of-the-art LLMs align with the moral values and hate speech labels annotated by our experts, we utilized the technique of LLMs as objective judges (Zheng et al., 2023; Verga et al., 2024; Ferron et al., 2023), and extended it with a multi-hop explanation evaluation strategy. This approach requires models not only to predict the labels of the input text, but also to infer the corresponding moral violation, with a brief evidence-based justification, specifying the rationales or group of words that triggered the moral value classification. We selected LLaMA-70B-Instruct [^4] https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct and GPT-4o Mini  [^5] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/, accessing them all via API. The prompting experiments were conducted on Google Colab using the CPU for processing, with the Macro F1 score metric from Sklearn library. LLaMA-70B-Instruct was chosen for its established effectiveness in similar evaluations Piot and Parapar (2025); Ngueajio et al. (2025) while GPT-4o-Mini provides an excellent balance of performance and cost-efficiency Nezhad and Agrawal (2025).
Moreover, using both models enables for a comparative analysis between an open (LLaMA) and closed-source model (GPT-4o-Mini). Parameter settings remained consistent across both models, with each being tasked with determining: (i) Whether the tweet comprises hate speech; (ii) Which moral sentiments are present in the tweet; (iii) Offering rationales for the moral sentiment prediction via the relevant text spans extracted from the tweet.
@@
To obtain comprehensive assessments, we implemented three distinct prompting strategies: (i) zero-shot prompting with no examples provided, asking the model to make determinations based solely on its knowledge; (ii) few-shot prompting with four examples included to help guide the model behavior; and Chain-of-Thought (CoT) strategy, with a step-by-step reasoning approach aimed at guiding the models’ reasoning. The examples of prompts are shown in the Appendix D. For each language, the four examples randomly selected for the few-shot prompting condition were removed from the datasets used in the zero-shot and CoT prompting conditions to allow a fair comparison.